Dry hands have a resistance of 1 Megaohm or more. If you happen to stick your hands in a live socket with 120 V you will feel a mild buzzing sensation in your hands, but you will remain unharmed. Why? Why is it that using a hairdryer is so hazardous in the shower, when you are wet? (Estimate the current in each case)
Answer #1

Ans

If the current though your hands(body) is

1) less than 5mA , then you feel mild buzzing sensation.

2) larger than 10mA can cause muscle spasms.

3) larger than 100mA can be very fatal.

The resistance of dry hands is about 1x106 ohms, so the current through your hands = 120/1x106 = 0.12mA i.e less than 5mA, hence mild buzzing sensation in your hands.

The resistance of wet skin is about 1x103 ohms, so current through your hands = 120/1x103 = 120mA which is greater than 100mA and can be very fatal.

So using hairdryer is hazardous in the shower as you are wet and current through you will be greater than 100mA.